### v2

# Quantum nonlinear effects in the number-conserving analogue gravity of Bose-Einstein condensates

Abstract
We consider the unitary quantum dynamics of a Bose-Einstein condensates at absolute zero, and demonstrate that an analogue gravity model in such a setting must take into account the backreaction of quasiparticle excitations onto the condensate background. This in turn requires that one expands to second order in perturbations, and takes the nonlinearity of the theory into account. It is shown that this leads to significant modifications of the standard linearized analogue gravity paradigm à la Unruh. In particular, to obtain a fully Lorentz-covariant equation in curved spacetime for second-order perturbations, we demonstrate that it is necessary to introduce, to leading order in powers of the formal expansion parameter $N^{-1/2}$ (where $N$ is total particle number), a quantum-fluctuation-renormalized spacetime metric which substantially differs from the Unruh metric and, to subleading order, two emergent vector fields. Both the renormalized metric and the vector fields then keep track of the backreaction of the quasiparticles onto the condensate up to the order in powers of $N^{-1/2}$ considered. Finally, we apply our formalism to an analogue-cosmological Friedmann-Lemaître-Robertson-Walker metric and establish its renormalized form due to the quantum many-body backreaction exerted by the excitation cloud.
I Introduction
The primary motivation behind the analogue gravity programme is the experimental simulation of fundamental features of field quantization in curved spacetimes [1] which are typically impossible to detect in real gravitational fields [2, 3, 4]. Starting with the work of Unruh [2], which put forth the idea that Hawking radiation can be observed in a lab fluid, analogue gravity models have now a rich history and found widespread applications to simulate various aspects of quantum field theory in curved spacetime cf., e.g., [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. A most recent review, containing numerous further references, is [16]; for a quick overview, a recent executive summary of analogue gravity’s major agenda has been provided in Ref. [17].
The analogue gravity idea of Unruh has led to the first observations of phenomena related to the paradigmatic quantum Hawking effect in the experimentally realizable physical system Bose-Einstein condensate (BEC), an ultracold quantum gas of interacting bosons [18, 19, 20], the analogue gravity arena we focus on in the present study. Other BEC analogue gravity experiments have probed aspects of cosmology, such as Hubble friction and preheating [21], as well as phenomena related to cosmological particle production [22, 23].
One crucial aspect of probing quantized fields in an arbitrary curved spacetime that is often neglected is the possible effect of the backreaction of the fluctuating field on the background itself. The relative magnitude of backreaction, say in terms of the contribution to the semiclassical energy-momentum tensor, must be much smaller than that of the gravitating matter field for a clear separation between what constitutes “background” and “fluctuation” to be maintained. To the lowest order, as in the original prediction of the Hawking effect, backreaction is indeed conventionally neglected [24], or at most a toy model for backreaction is proposed [25, 26]. However, a completely self-consistent treatment of any quantum field theory of a many-body system (or, similarly, of any candidate for an emergent unitary quantum field theory of gravity), should accommodate backreaction due to quantum-mechanical fluctuations created by its dynamical evolution. In this regard, analogue gravity models based on BECs represent an ideal testbed to explore the possible role of quantum backreaction, as these systems afford a fully controlled description in terms of Bogoliubov theory [27, 28, 29].
It is of major importance in what follows that for a proper treatment of backreaction problems in quantum systems such as BECs, as was first pointed out in [30], it is necessary to maintain particle-number conservation, cf. Refs. [31, 32, 33, 34], which are obtaining variants of number-conserving formulations of Bogoliubov theory which go beyond the usual symmetry-breaking, traditional Bogoliubov approach. To this end, in [30] the backreaction problem was defined by employing an expansion of the field operator in powers of $N^{1/2}$, where $N$ is the total number of particles (here either atoms or molecules) in the system. Then the backreaction of the quantum fluctuations can be elegantly treated as the subleading correction to the leading order in $N^{1/2}$ Gross-Pitaevskiǐ equation (GPE) [equivalent to a nonrelativistic variant of $\phi^{4}$ theory], incorporated by the modified coupled dynamical equation of the condensate and the non-condensate part. A particular lesson to be drawn from the analysis of [30] is that in a fully number-conserving description of the backreaction problem, the usual assumption of taking the expectation value of the quantized energy-momentum tensor, $\hat{T}_{\mu\nu}$, in a suitable quantum state, cf. Ref. [35], does not necessarily reproduce the correct, that is observable physical backreaction force.
The consequences of using a number-conserving approach were further explored in [36], where the solution for backreaction as a Cauchy problem was formulated, and solved for a dynamical BEC configuration. Importantly, it was shown in [36, 37] that there can exist a nonzero flux of condensate particles even in the absence of any particle flux due to quantum depletion, solely due to backreaction. Backreaction for a solitonic solution of the nonlinear Schrödinger equation was also studied most recently in [38].
In the present study, our aim is to build an analogue gravity model based on the number-conserving approach of [38], which incorporates the coupled dynamics (particle exchange) of the condensate and the depleted cloud. As we will explain in the sequel, the following is based on two key features that emerge from a consistent treatment of the backreaction of the depleted cloud onto the background [30, 39, 40, 36]. First of all, due to number conservation, there is a nonvanishing particle flux from the condensate to the depletion part. As a result, the continuity equations for density and current of both condensate and depletion contain drain (source) terms. In addition, the background to build an analogue gravity description now has a subleading contribution that represents a correction to the usual GPE. Consequently, as we argue, the standard approach of using strictly linear perturbation fields (see for example Visser’s pedagogical paper [3]) is not consistent with the order (in the perturbation expansion in $N$) of the field operator we will work with, ${\mathcal{O}}(N^{-1/2})$, which will be demonstrated to be the minimal expansion order necessary for number conservation to be maintained. We are thus required to go beyond the linear perturbation paradigm of Unruh. On the classical level, the latter was developed in Ref. [41], where it was shown how to extract an effective metric description from a generic nonlinear perturbation, at the cost of modifying the definition of background, also cf. Ref. [42].
We will argue that the above procedure engenders a significant departure from the established formulation of the fluid perturbation equations as those of a scalar field in curved spacetime and that it is not much less straightforward to conclude that such a description can be envisaged if both backreaction and nonlinearity are taken into account. We will indeed demonstrate that for a Lorentz-invariant form of the nonlinear perturbation equation for a scalar field to exist, the latter will acquire, in subleading order, a mass term and two “emergent” vector fields. To leading order in the large $N$ expansion, the effective curved background metric acquires corrections stemming from the exchange of particles between condensate and excitation cloud. As our analysis in this paper is based on the Bogoliubov approximation, i.e., the number of particles forming the condensate must be much larger than that of forming the depleted cloud, the backreaction corrections to the standard GPE are by necessity small (for the Bogoliubov theory to apply the gas must be dilute), and consequently also the above mentioned differences from the standard linearized analogue gravity procedure (which we dub in what follows the Unruh paradigm) must be small. However, due to the high tunability of quantum gases, in particular with respect to the two-body coupling strength, by Feshbach resonances (cf., e.g., Refs. [43, 44]), one can potentially observe such deviations from the conventional analogue gravity paradigm in a highly controlled manner in cold quantum gas experiments.
The paper is organized as follows: In Section. II, we explain the formulation of the number-conserving approach we follow throughout the paper. Section III consists of the description of identifying the Lorentz invariant dynamical equation for the perturbed field. Section IV describes the quantum nonlinear corrections to the lowest order semiclassical background metric and Section V elaborates on these corrections for a Friedmann-Lemaître-Robertson-Walker metric. In Section VI we conclude, and an Appendix contains the explicit evaluation of the drain term driving condensate loss for a Bose gas in a hard-walled box trap which is quenched from zero to finite interaction coupling.
II number-conserving formulation of the Bose-Einstein condensed gas
II.1 Governing field equations
The matter wave field $\Psi$ is assumed to be governed in the standard $s$-wave approximation appliicable to the dilute gas by the following field equation ($\hbar=1$)
$i\partial_{t}\Psi=\left(-\frac{1}{2}\nabla^{2}+V+U|\Psi|^{2}\right)\Psi,$ (1)
Here, $V$ is the external (trapping) potential, and $U>0$ is the interaction coupling from $s$-wave scattering between the gas particles. The theory described by (1) is invariant under global $U(1)$ transformations, which ensures the conservation of the total particle number $N$ according to the Noether theorem. The conservation law $\partial_{t}\rho+\nabla\cdot{\bm{J}}=0$ thus holds, where
$\rho=|\Psi|^{2},\qquad{\bm{J}}=\frac{1}{2i}(\Psi^{*}\nabla\Psi-\Psi\nabla\Psi^{*})$ (2)
are the system particle and current densities, respectively, and the total number of particles $N$ is given by the density integrated over all of space.
The variant of number-conserving theory we work with is based on the following formal expansion of the classical field in powers of $N^{1/2}$ [^1] Note there is a typo in Eq. (2) of Ref. [36], where ${\cal O}(N^{-3/2})$ should be replaced by $\mathcal{O}(N^{-1})$
$\Psi=\phi_{\rm 0}+\chi+\phi_{2}+\mathcal{O}(N^{-1}).$ (3)
Here, the scaling behavior of each components taking the limit of very large $N$ while keeping $UN=\mbox{constant}$, is given by [30]
$\phi_{0}=\mathcal{O}(N^{1/2}),\chi=\mathcal{O}(N^{0}),\ \mbox{and}\ \phi_{2}=\mathcal{O}(N^{-1/2}).$ (4)
We denote the quasiparticle (Bogoliubov) excitation field $\chi$ to keep the notation consistent with previous literature.
The above formal expansion (3) relies on the assumption that the number of non-condensed particles, $\delta N$, is much less than that of those forming the condensate part, and thus posits that $N_{0}\lesssim N$, also cf. Ref. [46] for the proof that in the mathematical limit $N\rightarrow\infty$ with $UN$ held fixed, the GPE (5) becomes exact. Here, $\chi$ is the fluctuation field associated with the non-condensed particles, and $\phi_{2}$ encompasses the correction of the condensate wave function due to the backreaction of the field $\chi$ onto the condensate. An expansion respecting the $N$ hierarchy is necessary to properly identify the fluctuation operator and the correction to the leading order Gross-Pitaevskiǐ orbital, see for details below.
The basic two equations determining the dynamics of the condensate and the noncondensate part in the full number-conserving formulations are the ${\cal O}{(N^{1/2})}$ GPE, which can be written as
$i\frac{\partial}{\partial t}\phi_{0}=\left[-\frac{1}{2}\nabla^{2}+V+U|\phi_{0}|^{2}\right]\phi_{0}.$ (5)
and the modified Bogoliubov–de Gennes (BdG) equation for the fluctuation field $\chi$
$i\frac{\partial}{\partial t}\chi=\left(-\frac{1}{2}\nabla^{2}+V+2U\rho_{0}\right)\chi+U\phi_{\rm 0}^{2}\chi^{*},$ (6)
with $\rho_{0}=|\phi_{0}|^{2}$. The above equation and the corresponding Hermitian conjugate equation are the modified versions of the Bogoliubov–de Gennes equations in the non-number-conserving standard Bogoliubov approach. When we consider the next-order corrections to the condensate wave function due to the backreaction of the non-condensate part of the system, as was done in [33, 34], then it can be shown that the $\mathcal{O}(N^{0})$ corrections to the leading order GPE are identically zero [33]. However in the next order ${\mathcal{O}}(N^{-1/2})$, the standard GPE is modified due to the correction from the time-dependent backreaction of the noncondensate particles, as described by the fluctuation operator. To provide a description of the corresponding mechanism, we will follow [30, 36] and define a quantum-fluctuation-corrected condensate wave function
$\phi_{c}=\phi_{0}+\phi_{2},$ (7)
which evolves according to the generalized GPE derived in Ref. [34],
$i\partial_{t}\phi_{c}=\Big{(}-\frac{1}{2}\nabla^{2}+V+U|\phi_{\rm c}|^{2}+2U|\chi|^{2}\Big{)}\phi_{c}+U\chi^{2}\phi_{c}^{*}.$ (8)
This equation describes the coupled dynamics of condensed and noncondensed parts taking the dynamical corrections due to the backreaction into account. We note that the corrected order parameter $\phi_{c}$ has a dominant ${\cal O}(N^{1/2})$ term plus a subdominant correction of order $N^{-1/2}$, and for a proper formulation of the backreaction, it is essential to take this correction into account.
To formulate the quantization of the fluctuation on top of the classical background constituted by $\phi_{c}$, we promote the fluctuation field $\chi$ and its Hermitian conjugate to be the fundamental quantum operators satisfying the standard bosonic commutation relations
$\left[\hat{\chi}({\bm{x}},t),\hat{\chi}^{\dagger}({\bm{x}}^{\prime},t)\right]=\delta({\bm{x}}-{\bm{x}}^{\prime}),$
$\left[\hat{\chi}({\bm{x}},t),\hat{\chi}({\bm{x}}^{\prime},t)\right]=\left[\hat{\chi}^{\dagger}({\bm{x}},t),\hat{\chi}^{\dagger}({\bm{x}}^{\prime},t)\right]=0.$ (9)
We will then assume that the instantaneous quantum state under consideration fulfills
$\langle\hat{\chi}({\bm{x}},t))\rangle\coloneqq 0\,\,\,\forall\,{\bm{x}},t\geq 0.$ (10)
This quantization choice, that is to regard $\hat{\chi}$ as the fundamental bosonic quantum field instead of $\hat{\Psi}$, does not maintain number conservation on the level of the full field operator of the bosonic particles: Due to $\langle\hat{\Psi}\rangle=\phi_{c}$, where however now $\phi_{c}$ is the backreaction-renormalized condensate mode in Eq. (7), the U(1) symmetry is still broken on the quantum level, as in the original Bogoliubov theory [27]. This choice, to promote $\chi$ to be the “fundamental” quantum field, however yields a set of equations amenable to a concrete solution the backreaction problem  [^2] We note in this connection that we were unable to locate an explicit backreaction calculation following the formally rigorous and in this regard very appealing approach of Ref. [33], which maintains $\langle\hat{\Psi}\rangle=0$ for any fixed $N$ state.. Furthermore, and most importantly in our context, it has been shown in [36] that with this vacuum choice for the quantized field $\hat{\chi}$, number conservation for condensate plus excitations is maintained, see Eqs. (18) and (19) below, and up to the order in $N^{-1/2}$ considered. The initial state for the Heisenberg picture evolution is then chosen as the noninteracting $U=0$ which has zero depletion $\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle=0\,\,(t=0)$ and $\phi_{2}=0$ (zero backreaction) [36] and can be set up in experiment in a well controlled manner. At $t=0$, we quench to a finite value of $U$ and follow the subsequent evolution.
The equation governing the dynamics of the corrected mean field is then given by the expectation value of (8)
$i\partial_{t}\phi_{c}=\left[-\frac{1}{2}\nabla^{2}+V+U|\phi_{\rm c}|^{2}+2U\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle\right]\phi_{c}+U\langle\hat{\chi}^{2}\rangle\phi_{c}^{*}.$ (11)
We will not delve into the details of defining the quasiparticle operator but will rather take the (pragmatic) viewpoint that field quantization can be performed at least for cases where the quasiparticle vacuum is well defined, such as in the simplest, that is noninteracting case. We therefore assume that we can evaluate terms such as the anomalous density appearing in the modified GPE by solving the $\mathcal{O}(N^{0})$ BdG equation as an explicit functional of the leading order GPE orbital $\phi_{0}$, along with the associated dependence on the spacetime coordinates $x,t$: see Appendix for a detailed calculation with the setup of Refs. [36, 37].
II.2 Quantum backreaction
We summarize here the formulation of the quantum backreaction problem in the formalism of Ref. [36], by taking the ${\mathcal{O}}(N^{-1/2})$ correction to the condensate background into consideration. The full system satisfies a conservation equation, $\partial_{t}\rho+\nabla\cdot{\bm{J}}=0$, where the average density is given as
$\rho:=\langle\hat{\rho}\rangle=\langle\hat{\Psi}^{\dagger}\hat{\Psi}\rangle,$ (12)
and the average current is
${\bm{J}}:=\langle\hat{\bm{J}}\rangle=\frac{1}{2i}\left\langle\hat{\Psi}^{\dagger}\nabla\hat{\Psi}-\nabla\hat{\Psi}^{\dagger}\hat{\Psi}\right\rangle.$ (13)
Utilizing the expansion of the full field operator (3) and evaluating the expressions above given the condition (10) we can write the density and current equations as
$\rho$ $=$ $|\phi_{c}({\bm{x}}|^{2}+\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle+{\mathcal{O}}(N^{-1/2}),$ (14)
${\bm{J}}$ $=$ $\frac{1}{2i}\Big{(}\phi^{*}_{c}\nabla\phi_{c}({\bm{x}}-\nabla\phi^{*}_{c}\Big{)}$ (15)
$+$ $\frac{1}{2i}\left\langle\hat{\chi}^{\dagger}\nabla\hat{\chi}-\nabla\hat{\chi}^{\dagger}\hat{\chi}\right\rangle+{\mathcal{O}}(N^{-1/2}).$
We have ordered the various terms by their order in $N^{1/2}$. For a proper description of the condensate correction due to quantum fluctuations it is useful to separate the contribution of the condensate and the noncondensate part in the full density and current as follows
$\rho_{\hat{\chi}}=\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle,$ (16)
${\bm{J}}_{\hat{\chi}}=\frac{1}{2i}\left\langle\hat{\chi}^{\dagger}\nabla\hat{\chi}-\nabla\hat{\chi}^{\dagger}\hat{\chi}\right\rangle,$ (17)
We then have $\rho=\rho_{c}+\rho_{\hat{\chi}}+{\mathcal{O}}(N^{-1/2})$ and $J=J_{c}+J_{\hat{\chi}}+{\mathcal{O}}(N^{-1/2})$.
Neither condensate nor depletion are separately conserved, even if only up to the order ${\mathcal{O}}(N^{0})$ we consider. In particular, we have for condensate and depletion cloud, respectively [36, 37],
$\partial_{t}\rho_{c}+\nabla\cdot{\bm{J}}_{c}$ $=$ $iU\Big{(}\phi^{2}_{c}\langle\hat{\chi}^{\dagger}\hat{\chi}^{\dagger}\rangle-\phi^{*2}_{c}\langle\hat{\chi}\hat{\chi}\rangle\Big{)}$ (18)
$\partial_{t}\rho_{\hat{\chi}}+\nabla\cdot{\bm{J}}_{\hat{\chi}}$ $=$ $-iU\Big{(}\phi^{2}_{c}\langle\hat{\chi}^{\dagger}\hat{\chi}^{\dagger}\rangle-\phi^{*2}_{c}\langle\hat{\chi}\hat{\chi}\rangle\Big{)}.$ (19)
We conclude that the theory is indeed conserving up to order ${\mathcal{O}}(N^{0})$ for density and current. This defines the extent to which our approach is number-conserving, as required. The number of particles in condensate and in the depletion cloud of quantum excitations are not separately conserved but only their respective sums, and thus the total density and current. The separate conservation equations for condensate and excitations acquire drain (source) terms which stem from backreaction.
III Analogue gravity in a number-conserving formalism
To understand how the propagation of fluctuations in a BEC is affected in a number-conserving description, we have to study how a perturbed solution of the equation (8) behaves for sufficiently small perturbation amplitude. We first decompose via the Madelung transformation the GPE orbital as
$\phi_{c}=\rho_{c}^{1/2}e^{i\theta_{c}},$ (20)
in terms of the two real functions $\rho_{c},\theta_{c}$. Then the modified GPE can be written in terms of the two real equations
$\partial_{t}\rho_{c}+\nabla\cdot\left(\rho_{c}\nabla\theta_{c}\right)=f_{1}(\rho_{c},\theta_{c},{\bm{x}},t),$ (21)
$\begin{split}\partial_{t}\theta_{c}+U\rho_{c}+\frac{1}{2}\left(\nabla\theta_{c}\right)\cdot\left(\nabla\theta_{c}\right)+f_{3}(\rho_{c},\theta_{c},{\bm{x}},t)\\ +V({\bm{x}},t)=0.\end{split}$ (22)
Here, we have defined the real functions $f_{1}(\rho_{c},\theta_{c},{\bm{x}},t)$, and $f_{3}(\rho_{c},\theta_{c},{\bm{x}},t)$, as
$f_{1}(\rho_{c},\theta_{c},{\bm{x}},t)$ $=$ $iU\Big{(}\phi^{2}_{c}\langle\hat{\chi}^{\dagger}\hat{\chi}^{\dagger}\rangle-\phi^{*2}_{c}\langle\hat{\chi}\hat{\chi}\rangle\Big{)},$
$f_{3}(\rho_{c},\theta_{c},{\bm{x}},t)$ $=$ $2U\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle+\frac{U}{2}\Big{(}\langle\hat{\chi}^{\dagger}\hat{\chi}^{\dagger}\rangle e^{2i\theta_{c}}+\langle\hat{\chi}\hat{\chi}\rangle e^{-2i\theta_{c}}\Big{)}.$
which are related to real and complex parts of the “anomalous” density. We have omitted the “quantum pressure” term $-h(\rho_{c})=-\frac{1}{2}\rho_{c}^{-1/2}\left(\nabla^{2}\rho_{c}^{1/2}\right)$ on the left-hand side of the Bernoulli Eq. (22); we will generally neglect density gradients in the background(s) in what follows. Including these gradients would lead us to the concept of (nonlocal) rainbow metrics [48]. The indices 1 and 3 on $f_{1}$ and $f_{3}$ indicate their respective orders ${\cal O}(N^{0})$ and ${\cal O}(N^{-1})$; see also the note after Eq. (26) below.
In the remainder of the paper, we will work with (21) and (22), which are essentially the standard fluid dynamical quasi-nonlinear equations, based on which we will explore our construction of analogue gravity in the second order nonlinear response regime. It is important to point out that the two functions $f_{1}$ and $f_{3}$ which contain the deviation from the usual hydrodynamic description based on the usual leading order GPE are of the order $\mathcal{O}(N^{0})$ and $\mathcal{O}(N^{-1})$, respectively. Thus they are subdominant contributions to Eqs. (21 and 22). They however encapsulate the backreaction we are interested in, and thus we fully keep $f_{1},f_{3}$, assuming them to be general functions of $\rho_{c}$, and $\theta_{c}$. Note that $f_{1}<0$ corresponds to the loss of particles from the condensate.
Let us now elaborate on two important differences from conventional analogue gravity models obtained from the hydrodynamical equations by using the standard GPE (5):
(1) If we incorporate the action of the noncondensed part in the evolution, then the condensed part of the continuity equation for $\phi_{c}$ is not separately conserved, and this is embedded in the drain term $f_{1}(\rho_{c},\theta_{c})$. As a result, the standard continuity equation is no longer valid, and we have to incorporate the breakdown of continuity in the formulation of the analogue gravity description based on this background.
(2) Once we include the backreaction field $\phi_{2}$, then the background on which we aim to build an analogue gravity model, namely $\phi_{c}$ has two pieces, one is the usual dominant ${\mathcal{O}}(\sqrt{N})$ contribution $\phi_{0}$ and the other is the subdominant $\phi_{2}$, which is of order ${\mathcal{O}}(N^{-1/2})$. To encapsulate the effect of backreaction in the analogue gravity description, the perturbation expansion of the solution of the generalized GPE around a sufficiently slowly varying background should thus also be of ${\mathcal{O}}(N^{-1/2})$. It is not sufficient just to consider linearized perturbations ${\mathcal{O}}(N^{0})$ on top of a background including backreaction, as the background itself contains the field $\phi_{2}={\mathcal{O}}(N^{-1/2})$.
In summary, we have to go beyond the standard linear fluctuation analysis to see how the backreaction of the quasiparticle excitation field changes the standard analogue gravity description [41].
III.1 Linearized fluctuations around a background
We first briefly review the basic structure of the conventional analogue gravity formulation based on linear fluctuations around a background [2, 3, 4].
The fundamental two equations describing the evolution of the fluid variables $\rho_{c}$, and $\theta_{c}$, namely the (modified) continuity equation (21) and the Euler equation (22) are coupled first-order quasilinear partial differential equations. They have a unique solution for well-defined boundary conditions. The choice of “background” and “perturbation” is always to some extent arbitrary. Conventionally one however chooses the spatiotemporally slowly varying part of the full evolution to constitute the background, which then also justifies our neglect of background density gradient terms throughout.
We will assume a convergent power series expansion of the fluid variables ($\rho_{c}$, $\theta_{c}$) around the background ($\rho_{0}$, $\theta_{0}$), in the form [49]
$\rho_{c}=\rho_{0}+\epsilon\delta\rho_{1}+\epsilon^{2}\delta\rho_{2}+...,$ (24)
$\theta_{c}=\theta_{0}+\epsilon\delta\theta_{1}+\epsilon^{2}\delta\theta_{2}+.....$ (25)
Here, the scaling of the density contributions with the number of particles is as follows,
$\rho_{0}=\mathcal{O}(N^{1/2}),\quad\delta\rho_{1}=\mathcal{O}(N^{0}),\quad\delta\rho_{2}=\mathcal{O}(N^{-1/2})$ (26)
and so on. For later convenience, we can thus establish a one-to-one correspondence with the subscript of each component of the density perturbation with the relevant scaling with $N^{1/2}$. For example, the subscript $0\rightarrow\mathcal{O}(N^{1/2})$ and similarly for other terms in descending order, i.e., an index $\alpha$ implies being of ${\cal O}(N^{(1-\alpha)/2})$. Note also that $f_{1}=\mathcal{O}(N^{0})$ and $f_{3}=\mathcal{O}(N^{-1})$ defined earlier in (LABEL:f1f2def) follow the same scaling convention, but that the phase perturbation index in (25) does not indicate a scaling with powers of $N^{1/2}$.
We stress here for clarity that we are dealing with two expansion parameters in terms of small quantities, where $N^{-1/2}$ is the mean-field expansion parameter and $\epsilon$ is the parameter indicating the amplitude perturbation order of the fields. If we choose to follow the Unruh paradigm [4, 2], then it would suffice to keep only ${\mathcal{O}}(\epsilon)$ terms, and we can study the dynamics of the perturbed phase $\delta\theta$ around the background $\rho_{0},\theta_{0}$, using that the continuity equation remains satisfied. Then it is possible to describe the dynamics of the (first-order) fluctuation field $\delta\theta_{1}$ as a massless Klein-Gordon equation in an effective curved spacetime
$\frac{1}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\delta\theta_{1}\right)=0.$ (27)
The components of the metric $g_{\mu\nu}$ depend on the background variables $\rho_{0},\theta_{0}$, where the speed of sound in this background is $c^{2}_{s0}=U\rho_{0}$.
Importantly, this procedure implies that we are taking the background variables $\rho_{0},\theta_{0}$, and the perturbed variables $\rho_{1}=\rho_{0}+\epsilon\delta\rho_{1},\theta_{1}=\theta_{0}+\epsilon\delta\theta_{1}$ to be two independent sets of solutions of the fluid-dynamical equations, for proper boundary conditions. It is, doing so, assumed that the boundary conditions are such that expansions of the form (24) and (25) are possible, i.e, that the corresponding series do not diverge.
III.2 Second-order fluctuations around a background
As explained in the previous section, the orbital $\phi_{c}$ has also a subleading contribution, and if we want to see how the effect of such a subleading contribution changes the conventional Unruh paradigm, then it is essential to consider the series expansion (24) at least up to $\delta\rho_{2}$. However, it is also important to understand how the terms $\mathcal{O}(N^{0})$, i.e. $\delta\rho_{1}$, change the previous background $\rho_{0},\theta_{0}$, and if it is at all possible to provide an analogue gravity description with respect to the new background [41].
Let us first assume perturbation amplitude expansions as in Eqs. (24) and (25)) are possible, up to the indicated quadratic order. Then from the modified continuity equation (21), and the Euler equation 22, we obtain the following set of equations:
$\partial_{t}(\rho_{0}+\epsilon\delta\rho_{1})+\nabla\cdot\Big{(}(\rho_{0}+\epsilon\delta\rho_{1})(\nabla\theta_{0}+\epsilon\nabla\delta\theta_{1})\Big{)}=f_{1}(\rho_{0},\theta_{0})+\epsilon\left.\frac{\partial f_{1}}{\partial\rho_{c}}\right|_{0}\delta\rho_{1}+\epsilon\left.\frac{\partial f_{1}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}+\epsilon^{2}\left.\frac{\partial^{2}f_{1}}{\partial\rho_{c}\partial\theta_{c}}\right|_{0}\delta\rho_{1}\delta\theta_{1},$
$\partial_{t}(\theta_{0}+\epsilon\delta\theta_{1})+U(\rho_{0}+\epsilon\delta\rho_{1})+\frac{1}{2}(\nabla(\theta_{0}+\epsilon\delta\theta_{1}))^{2}+V+f_{3}(\rho_{0},\theta_{0})+\epsilon\left.\frac{\partial f_{3}}{\partial\rho_{c}}\right|_{0}\delta\rho_{1}+\epsilon\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}+\epsilon^{2}\left.\frac{\partial^{2}f_{3}}{\partial\rho_{c}\partial\theta_{c}}\right|_{0}\delta\rho_{1}\delta\theta_{1}=0.$
where we used the abbreviation $\left.\cdots\right|_{0}\coloneqq\left.\cdots\right|_{\rho_{0},\theta_{0}}$ here and in what follows. As we will argue later in detail, the two above equations define the new background on which we will consider the propagation of nonlinear fluctuations. Then the new equations governing the perturbations are
$\epsilon^{2}\partial_{t}\delta\rho_{2}+\nabla\cdot\Big{(}\epsilon^{2}\delta\rho_{2}(\nabla\theta_{0}+\epsilon\nabla\delta\theta_{1})+\epsilon^{2}(\rho_{0}+\epsilon\delta\rho_{1})\nabla\delta\theta_{2}\Big{)}=\epsilon^{2}\left.\frac{\partial^{2}f_{1}}{\partial\rho_{c}^{2}}\right|_{0}\delta\rho_{2}+\epsilon^{2}\left.\frac{\partial^{2}f_{1}}{\partial\theta_{c}^{2}}\right|_{0}\delta\theta_{2},$ (30)
$\epsilon^{2}\partial_{t}\delta\theta_{2}+\epsilon^{2}U\delta\rho_{2}+\epsilon^{2}\nabla(\theta_{0}+\epsilon\delta\theta_{1})\cdot\nabla\delta\theta_{2}+\epsilon^{2}\left.\frac{\partial^{2}f_{3}}{\partial\rho_{c}^{2}}\right|_{0}\delta\rho_{2}+\epsilon^{2}\left.\frac{\partial^{2}f_{3}}{\partial\theta_{c}^{2}}\right|_{0}\delta\theta_{2}=0.$ (31)
First of all, here we have, in accordance with the subleading contribution to the background including backreaction, used an inhomogeneous separation of effective contributions, controlled by the separation parameter $\epsilon$ (inhomogeneity in this parameter), to transform the initial fluid-dynamical equations for the fields into a set of four equations for “background,” Eqs. (28) and (29) and “perturbation,” (30) and (31), cf.  the corresponding procedures in Ref. [50]. As we will see, this is mandatory to obtain an analogue gravity description of these nonlinear perturbation equations.
Let us consider the following combinations of variables,
$\rho_{b}=\rho_{0}+\epsilon\delta\rho_{1},\quad\theta_{b}=\theta_{0}+\epsilon\delta\theta_{1},$ (32)
${\rho_{c}}=\rho_{0}+\epsilon\delta\rho_{1}+\epsilon^{2}\delta\rho_{2},\quad{\theta_{c}}=\theta_{0}+\epsilon\delta\theta_{1}+\epsilon^{2}\delta\theta_{2},$ (33)
where we truncated in the second line the series in Eqs. (24) and (25) after second order in $\epsilon$. We assume that these combinations satisfy the dynamical fluid Eqs. (21) and (22)) for two different boundary conditions, such that the expansions above are convergent, which in turn requires $\epsilon^{2}\delta\rho_{2}\ll\epsilon\delta\rho_{1}\ll\rho_{0}$. This also implies that we are not separately demanding $\rho_{0},\theta_{0}$ to follow any dynamical equations, at least formally. Then the condition on $\rho_{0},\theta_{0}$ is simply that they are two sufficiently slowly varying “background” functions, which are not necessarily solutions to the fluid dynamical equations, but can be chosen, for example, according to the ease of interpretation of experiments. Namely, such that the first and second order perturbative corrections on top of the background satisfy the fluid-dynamical equations.
We can consider of course also fluctuations higher than of quadratic order and, in a similar fashion, demand that we can suitably define a background for the next-order density and phase fluctuations. In a minimally self-consistent analogue gravity model, as a consequence of the backreaction due to enforcing number conservation, the perturbed fluid variables have however to be at least of quadratic order, i.e., should scale here as $\mathcal{O}(N^{-1/2})$, to be properly treated as a perturbation.
With respect to the variables ($\rho_{b},\theta_{b}$) of Eq. (32) defining the new background, we can substitute $\delta\rho_{2}$ from (31) into (30) to obtain
$\partial_{t}\Big{(}\frac{1}{\tilde{U}}\Big{[}\partial_{t}\delta\theta_{2}+{\bm{v}}_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{]}\Big{)}+\nabla\cdot\Bigg{(}\frac{1}{\tilde{U}}{\bm{v}}_{b}\Big{[}\partial_{t}\delta\theta_{2}+{\bm{v}}_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{]}-\rho_{b}\nabla\delta\theta_{2}\Bigg{)}\\ -\frac{\ddot{f}_{1}}{\tilde{U}}\Big{(}\partial_{t}\delta\theta_{2}+{\bm{v}}_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{)}+f^{\prime\prime}_{1}\delta\theta_{2}=0.$ (34)
Here, we have indicated partial derivatives with respect to $\rho_{c}$, and $\theta_{c}$ as overdot and prime, respectively. Furthermore, we have defined the background velocity
${\bm{v}}_{b}=\nabla(\theta_{0}+\epsilon\delta\theta_{1}),$ (35)
and an effective coupling constant
$\tilde{U}=U+\left.\frac{\partial^{2}f_{3}}{\partial\rho_{c}^{2}}\right|_{0}.$ (36)
Note that $U$ and $f_{3}$ both scale as ${\cal O}(N^{-1})$.
III.3 Lorentz-invariant wave equation for second-order perturbations
As the continuity equations are not satisfied separately, e.g., the condensate loses particles, there are extra terms present that indicate that the dynamics of the perturbation field $\delta\theta_{2}$ can not be written simply as a Lorentz-invariant curved spacetime equation.
However, it is possible to obtain a Lorentz invariant wave equation for the dynamics of the perturbation field. This comes at the price of introducing additional fields, though, as we will now discuss in detail. We obtain from Eq. (34) (without any further approximation) the following equation of motion
$\frac{1}{\sqrt{-\mathbf{g}}}\mathcal{D}_{\mu}\left(\sqrt{-\mathbf{g}}\mathbf{g}^{\mu\nu}\mathcal{D}_{\nu}\delta\theta_{2}\right)\\ -\mathbf{g}^{\mu\nu}\mathcal{A}_{\mu}\mathcal{D}_{\nu}\delta\theta_{2}-\mathcal{B}^{\mu}\mathcal{D}_{\mu}\delta\theta_{2}+\mathcal{M}^{2}\delta\theta_{2}=0.$ (37)
Here, we have defined the modified derivative operator as
$\mathcal{D}_{\mu}=\partial_{\mu}+\mathcal{A}_{\mu}.$ (38)
Note that $A_{\mu}$ does not assume the role of a gauge field here and is defined such that it is real; see Eq. (42) below.
We can identify the analogue metric components as follows
$\mathbf{g}_{tt}=-\sqrt{\frac{\rho_{b}}{\tilde{U}}}\left(\tilde{U}\rho_{b}-{\bm{v}}_{b}^{2}\right),$ (39)
$\mathbf{g}_{ti}=-\sqrt{\frac{\rho_{b}}{\tilde{U}}}{\bm{v}}_{bi},\qquad\mathbf{g}_{ij}=\sqrt{\frac{\rho_{b}}{\tilde{U}}}\delta_{ij},$ (40)
where the metric determinant is given by
$\sqrt{-\mathbf{g}}=\sqrt{\frac{\rho_{b}^{3}}{\tilde{U}}}.$ (41)
Here, the index $i$ represents the spatial coordinates.
We have defined in Eqs. (37) and (38) the components of a vector field
$\mathcal{A}^{\mu}=\Big{(}\frac{f^{\prime\prime}_{3}}{\sqrt{\mathbf{-g}}\tilde{U}},\frac{f^{\prime\prime}_{3}{\bm{v}}_{b}}{\sqrt{\mathbf{-g}}\tilde{U}}\Big{)},$ (42)
and the components of a second vector field $\mathcal{B}^{\mu}$ as
$\mathcal{B}^{\mu}=\Big{(}\frac{\ddot{f}_{1}}{\sqrt{\mathbf{-g}}\tilde{U}},\frac{\ddot{f}_{1}{\bm{v}}_{b}}{\sqrt{\mathbf{-g}}\tilde{U}}\Big{)}.$ (43)
Finally, the effective mass of the scalar field is identified as follows
$\mathcal{M}^{2}=\frac{1}{\sqrt{\mathbf{-g}}}\Big{[}-\frac{2\ddot{f}_{1}f^{\prime\prime}_{3}}{\tilde{U}}+f^{\prime\prime}_{1}\Big{]}.$ (44)
Note that spacetime indices are raised and lowered throughout with the full metric $\mathbf{g}_{\mu\nu}$, not with the Unruh metric $g_{\mu\nu}$.
As can be readily concluded from the above Eqs. (42, (43), and (44)), the corresponding fields are identically zero when the two functions $f_{1},f_{3}$ vanish, producing the result of the Unruh paradigm in terms of an effective metric only then also to second order in perturbations.
In summary, the dynamics of the nonlinear phase fluctuations can be written, to second order in the perturbations, as a generalized massive Klein-Gordon equation in a curved spacetime, coupled with two emergent vector fields, all with respect to the background $\rho_{b},{\bm{v}}_{b}$.
The contributions of the final two (vector field) terms in the dynamics of the scalar field are subdominant (are of lower order in $N^{-1/2}$) when compared to the first two terms. The second-order emergent metric $\mathbf{g}_{\mu\nu}$ includes a term $h_{\mu\nu}$ which signifies the deviation from the Unruh paradigm metric $g_{\mu\nu}$. The $N$-scalings of the various terms are as follows
$\begin{split}\mathbf{g}_{\mu\nu}=g_{\mu\nu}[\mathcal{O}(N)]+h_{\mu\nu}[\mathcal{O}(N^{0})],\\ A_{\mu}=[\mathcal{O}(N^{-1})],\\ B_{\mu}=[\mathcal{O}(N^{-2})].\end{split}$ (45)
where $g_{\mu\nu}[\mathcal{O}(N)]$ is the conventional Unruh paradigm metric. Hence to leading order in the large $N$ expansion, corrections to the Unruh metric dominate the quantum nonlinear effects in the Lorentz-invariant wave equation (37).
In our context of the BEC-based analogue gravity model incorporating number conservation, the above corrections to the Unruh paradigm can be readily understood by considering a system where the interaction is turned on suddenly in an initially noninteracting gas [36], also see the Appendix. Then, at sufficiently early time, when the number of depleted particles is very small, the correction due to backreaction is also negligible, and the Unruh paradigm metric is an excellent description. However, as the system evolves, the number of depleted particles is no longer negligible, and the corrected metric will afford a better description.
The equation (37) is not invariant under time reversal $t\rightarrow-t$. The presence of a first-order derivative in time breaks the time reversal symmetry, which in turn can be traced back to the existence of particle exchange between the condensed and the noncondensed parts of the system, which is fully respected in our construction of a number-conserving theory. The manifestation of the breaking of time reversal is a complex quasiparticle frequency, cf. Eq. (47) below.
III.4 Comparison to other approaches
At this point, it is appropriate to draw a comparison between the formulation we have adopted here and in particular the one proposed in the classical analysis of [41], where the propagation of a general nonlinear wave in a fluid medium described by the coupled continuity and Euler equations was considered from the viewpoint of analogue gravity.
It was shown in [41] that an effective metric can be derived also when the intrinsic nonlinearity of a (classical perfect) fluid is taken into account. This description however breaks down after a certain time scale due to the inherent nonlinear nature of the fluid motion and the consequent creation of shocks. It is nevertheless possible to obtain a spacetime metric description provided we take the backreaction of the nonlinear motion into account and consequently modify the background itself. In the present paper, we have taken a weaker restriction on the nature of nonlinearity in the fluid under consideration. To be more specific, distinct from [41], we have assumed that the perturbation expansion in (24) is convergent, i.e., the successive terms in this expansion are gradually of sufficiently lower order than the previous one, an assumption that was not a priori taken in [41]. Also our analysis will break down after a certain timescale at which the fully nonperturbative nature of the wave propagation can no longer be neglected. The exact relation between this timescale, and the failure of the Bogoliubov approximation, i.e., when the number of condensate particles becomes of order the number of noncondensed particles, should be an interesting topic to pursue in the future.
Approach Background considered Number-conserving References
Leading-order GPE (5)    Linearized on top of $(\rho_{0},{\bm{v}}_{0})$ No [8, 9, 12]
Number-conserving ansatz (3) Full fluid density and velocity [^3] In Refs. [30],and [36] the importance of full number conservation was pointed out. However, no consequences for the analogue metric      to any given order in the large $N$ expansion were considered. Yes [30]
Nonlinear waves in classical fluid Nonperturbative fluctuation (redefined background) [^4] The formulation in [41] is for a generic barotropic, non-viscous, and irrotational fluid, with a BEC as a particular case. However, the redefined background in [41] is not the $(\rho_{b},{\bm{v}}_{b})$ employed here. N/A [41]
Corrected GPE (11) Second-order on background $(\rho_{b},{\bm{v}}_{b})$ Yes [36], This work
Table 1: A table comparing formulations used to construct analogue gravity models for quantum and classical fluids.
In the context of the present work, as there is a well defined separation in powers of $N^{-1/2}$, it is essential to explore in detail how the dynamics of nonlinear perturbations changes due to backreaction. In fact, as the backreaction of the noncondensed part is $\mathcal{O}(N^{-1/2})$, a fully nonlinear analysis as in [41] might not be sensitive (in a controlled manner) to this correction.
We have summarized the features of various approaches to analogue gravity in quantum and classical fluids which have been taken since its inauguration, in table 1, together with a few selected references. The combination of generality and degree of control possible within our quantum many-body approach (which can be extended to higher orders of powers of $N^{-1/2}$) becomes apparent. Most importantly, we provide the first spacetime metric description of the propagation of second-order perturbations in a quantum fluid, within a number-conserving formulation of backreaction.
III.5 Dispersion relation
To understand the nature of the dispersion relation, let us consider a plane wave propagating in $x$ direction (37)
$\delta\theta_{2}\propto\exp[i(kx-\omega t)],$ (46)
where the frequency $\omega$ can in general be complex. Substituting this ansatz back into Eq. (34), and assuming that spatiotemporal variations of the background are negligible, we have
$\omega=v_{bx}k+\frac{1}{2}\left[i(\ddot{f}_{1}-f^{\prime\prime}_{3})\right.\\ \left.\pm\sqrt{4(\tilde{U}f^{\prime\prime}_{1}-\ddot{f}_{1}f^{\prime\prime}_{3}+c^{2}_{s,b}k^{2})-(\ddot{f}_{1}-f^{\prime\prime}_{3})^{2}}\right].$ (47)
Here, we have defined the effective sound velocity of the background ($\rho_{b}$, $\theta_{b}$) to be $c_{s,b}^{2}=\tilde{U}\rho_{b}$. As expected, when $f_{1}=f_{3}=0$, this dispersion relation reduces to the standard Bogoliubov dispersion. The appearance of a generally complex frequency is a direct manifestation of backreaction, causing the condensate particle number to be not separately conserved. The exact rate at which this occurs, and the corresponding correction on the leading-order Bogoliubov dispersion depends on the relevant field configurations leading to the solutions of the BdG equation (6).
In the Appendix, we discuss a specific time dependent configuration for a BEC contained in a box potential, where the quantization of the depletion field $\hat{\chi}$ can be performed analytically [36]. Consequently, it is possible to calculate the anomalous correlators appearing in the definitions of $f_{1}$ and $f_{3}$. As we explain in the Appendix in detail, it can be concluded that for early times, after switching on interactions from an initially noninteracting state, the particle loss from the condensate, quantified by the function $f_{1}$, is quadratic in time, and this loss is in such a setup generic in the sense that it is independent of the system size, at least for sufficiently early times. As a result, the leading order correction to the Unruh paradigm metric is also generic and governed by $h_{\mu\nu}$ in (45), until the Bogoliubov approximation itself breaks down.
IV Quantum nonlinear spacetime metric
Even though the form of the analogue metric that couples to the second-order fluctuation field, $\mathbf{g}_{\mu\nu}$ is formally identical to that of a metric corresponding to the linear regime, the exact functional dependence on the underlying fluid variables is different, due to the fact that now the definition of background has changed. In this section we will explicitly calculate the components of the metric and will see how they differ from the standard Unruh paradigm. In the remainder of the paper we will denote the emergent nonlinear background as background-II (variables $\rho_{b}$, $\theta_{b}$), while the background related to the strict linearization leading to the Unruh paradigm will be denoted without italics and boldface as background-I (variables $\rho_{0}$, $\theta_{0}$). For both backgrounds we demand, as usual, that only perturbations which vary much faster than the background are considered. In particular, the relevant length scales of background-II must always be much less than those of background-I for them to be meaningfully separable.
IV.1 Metric coupled to second-order fluctuations
Expanding the new metric components given by Eqs. (39) and (40), and using the fluid variables $\rho_{b},\theta_{b}$ to define the new background-II we obtain
$\mathbf{g}_{tt}=g_{tt}-\sqrt{\frac{\rho_{0}}{U}}\Big{[}\epsilon U\delta\rho_{1}-2\epsilon{\bm{v}}_{0}\cdot{\bm{v}}_{1}-\epsilon^{2}{\bm{v}_{1}}^{2}\\ +\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\big{(}c_{s0}^{2}-{\bm{v}}^{2}_{0}+\epsilon U\delta\rho_{1}-2\epsilon{\bm{v}}_{0}\cdot{\bm{v}}_{1}\big{)}-\epsilon^{2}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}\big{(}c_{s0}^{2}-{\bm{v}}^{2}_{0}\big{)}\Big{]},$ (48)
$\mathbf{g}_{ti}=g_{ti}-\sqrt{\frac{\rho_{0}}{U}}\Big{[}\epsilon{\bm{v}}_{1i}+\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\big{(}{\bm{v}}_{0}+\epsilon{\bm{v}}_{1}\big{)}_{i}-\epsilon^{2}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}{\bm{v}}_{0i}\Big{]}.$ (49)
$\mathbf{g}_{ij}=g_{ij}+\sqrt{\frac{\rho_{0}}{U}}\Big{[}\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}.$ (50)
Here, we have defined the first-order correction to the velocity field as ${\bm{v}}_{1}=\nabla\delta\theta_{1}$, and retained terms up to $\mathcal{O}(\epsilon^{2})$ in the first-order perturbed fields $\delta\rho_{1}$ and $\delta\theta_{1}$, which represents the working order for the new background-II.
The standard linearized Unruh paradigm metric has components
$g_{tt}=-\sqrt{\frac{\rho_{0}}{U}}\left(c_{s0}^{2}-{\bm{v}}_{0}^{2}\right),\qquad g_{ti}=-\sqrt{\frac{\rho_{0}}{U}}{\bm{v}}_{0i},\\ g_{ij}=\sqrt{\frac{\rho_{0}}{U}}\delta_{ij}.$ (51)
The fact that the leading-order contribution to the effective metric that describes the dynamics of the second-order fluctuation field $\delta\theta_{2}$ is given by the Unruh metric components related to background-I indicates that the spacetime metrics corresponding to the two backgrounds are related by
$\mathbf{g}_{\mu\nu}=g_{\mu\nu}+h_{\mu\nu},$ (52)
where $h_{\mu\nu}$ can be thought of as emergent, as it is due to the inherent nonlinearity of background-II. From Eq. (45), we conclude that the leading-order corrections come from the nonlinear contribution to the modified metric ($h_{\mu\nu}$) with respect to the background-II. In fact, as might be expected on general physical grounds, this correction is of the same order as the density of the depletion cloud $\rho_{\hat{\chi}}=[\mathcal{O}(N^{0})]$.
IV.2 Dynamics of the background field generating curved spacetime
To understand how changing the form of background affects the analogue gravity description in our formalism, we start with a simple example. Let us consider a uniform and static fluid medium in one spatial dimension, described by the variables $\rho_{0},\theta_{0}$. Then the metric for linear sound propagation is Minkowski, which corresponds to background-I. Now we consider the propagation of a nonlinear perturbation; then, as we described in the above, it is possible to obtain a Lorentz-invariant description only when we couple the second-order field $\delta\theta_{2}$ with the background-II, which now however does not represent a simple Minkowski spacetime. The components of the tensor $h_{\mu\nu}$ in this case can be obtained by setting ${\bm{v}}_{0}=0$ and $\rho_{0}=$ constant in Eqs. (48), (50), and (49). It is also important to note that we have imposed an extra condition on the fluid variables, namely that the initial configuration $\rho_{0},\theta_{0}$ satisfies the fluid-dynamical equations separately, which however is not necessary to obtain the metric for second-order fluctuation fields. This relates to employing distinct boundary and initial conditions, which are imposed on the quasi-linear fluid dynamical equations we are dealing with, and again emphasizes how the distinction between the background and the fluctuation fields is to some extent arbitrary and can thus also change an analogue gravity description significantly.
The dynamical equation satisfied by this scalar field generating curved spacetime can be obtained from the two equations that define background-II. Substituting $\delta\rho_{1}$ from Eq. (29) into Eq. (28), we obtain the nonlinear equation of motion for the field $\delta\theta_{1}$ as, cf. Ref. [41],
$\begin{split}\partial_{t}\Big{[}-\frac{1}{\lambda}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\Big{]}+\nabla\cdot\Big{[}\rho_{0}\nabla\delta\theta_{1}-\frac{{\bm{v}}_{0}}{\lambda}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\\ -\frac{\epsilon}{\lambda}(\nabla\delta\theta_{1})\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\Big{]}\\ +\frac{1}{\lambda}\left.\frac{\partial f_{1}}{\partial\rho_{c}}\right|_{0}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}-\left.\frac{\partial f_{1}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\\ -\frac{\epsilon}{\lambda}\left.\frac{\partial^{2}f_{1}}{\partial\rho_{c}\partial\theta_{c}}\right|_{0}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\delta\theta_{1}=0.\end{split}\vspace{-6em}$ (53)
We have defined the first-order modification to $U$ as
$\lambda=U+\left.\frac{\partial f_{3}}{\partial\rho_{c}}\right|_{0}$ (54)
neglecting a sub-leading correction of $\mathcal{O}(\epsilon^{2})$, which should be compared with its second-order counterpart in Eq. (36).
The above highly nonlinear equation (53) describes the dynamics of the spacetime background-II due to the intrinsically nonlinear nature of quantum fluctuations. We apply it in the next section to a cosmological scenario.
V Quantum nonlinear corrections to a cosmological background
V.1 Friedmann-Lemaître-Robertson-Walker background
We will now discuss in more detail how the quantum nonlinear effects coming from the backreaction of the quantized field $\hat{\chi}$ change the leading-order Unruh paradigm metric, which governs perturbations to linear order. To exemplify our arguments in the preceding sections we will consider a solution of the GPE for $\phi_{0}$, such that the leading-order $\mathcal{O}(N)$ density $\rho_{0}$ is constant and the condensate is at rest. This produces a simple analogue of the Friedmann-Lemaître-Robertson-Walker (FLRW) metric of cosmology when $c_{s0}=c_{s0}(t)$, cf. Ref. [51], where the time dependence of the sound speed is implemented by using a Feshbach resonance sweep [44] such that $U=U(t)$, with the constant density enforced by a hard-walled trap of infinite height.
Then the Unruh paradigm metric, as seen by the first order fluctuation field $\delta\rho_{1},\delta\theta_{1}$, is given by
$g^{\rm FLRW}_{\mu\nu}=\sqrt{\frac{\rho_{0}}{U(t)}}\begin{bmatrix}-{c_{s0}^{2}(t)}&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix},$ (55)
where the spatial geometry is conformally flat, according to the usual assumption that the (analogue) universe is isotropic and homogeneous at the cosmological scale. In this model, a decreasing $U(t)$ will produce a spacetime corresponding to an expanding universe, and vice versa. The metric of (55) refers to the background-I, which consists of fluid variables $\rho_{0},\theta_{0}$, that is, solutions of the leading-order GPE (5).
However, as we explained in sections III and IV, the spacetime metric as seen by the second-order fluctuation field $\delta\theta_{2}$ is not the simple FLRW metric of (55), as the field $\delta\theta_{2}$ does not couple to the background-I, rather it couples to the redefined background-II defined in terms of the fluid variables $\rho_{b}$, $\theta_{b}$. As a result, the leading-order FLRW type Unruh metric will receive, in that perturbative order, subleading corrections of $\mathcal{O}(N^{0})$, the general form of which is quoted in Eqs. (48), (49), and (50), and the dynamics of which is governed by (53).
In the next subsection, we study the form of the quantum fluctuation renormalized metric $\mathbf{g}_{\mu\nu}$ when the first-order fluctuation field $\delta\theta_{1}$ perceives a Minkowski or FLRW Unruh paradigm metric $g_{\mu\nu}$.
V.2 Renormalized FLRW Unruh metric
As stated, the background-I ($\rho_{0}$, $\theta_{0}$), corresponding to the first order fluctuation field ($\delta$$\theta_{1}$) is a FLRW metric if the sound speed in BEC is time dependent. On the other hand, the metric as seen by the second-order fluctuation field $\delta\theta_{2}$, which only couples with the background-II, a solution of the corrected GP equation (11), is not the simple FLRW metric of the form (55). The latter changes due to backreaction, from Eqs. (48)–(50), as follows
$\mathbf{g}_{tt}=g^{\rm FLRW}_{tt}-\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}\Big{(}\epsilon U(t)\delta\rho_{1}-\epsilon^{2}{\bm{v}}^{2}_{1}\Big{)}\\ +\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\Big{(}U(t)\rho_{0}+\epsilon U(t)\delta\rho_{1}\Big{)}-\epsilon^{2}U(t)\rho_{0}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}\Big{]}\\ =-\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}U(t)\rho_{0}+\frac{3}{2}\epsilon U(t)\delta\rho_{1}+\frac{3}{8}\epsilon^{2}U(t)\frac{\delta\rho^{2}_{1}}{8\rho_{0}}\Big{]},$ (56)
$\mathbf{g}_{ti}$ $=$ $g^{\rm FLRW}_{ti}-\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}\epsilon{\bm{v}}_{1i}+\epsilon^{2}\frac{\delta\rho_{1}}{2\rho_{0}}{\bm{v}}_{1i}\Big{]},$ (57)
$\mathbf{g}_{ij}$ $=$ $g^{\rm FLRW}_{ij}+\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}$ (58)
$=$ $\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}1+\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}.$
Here the boundary conditions are chosen such that the background-I ($\rho_{0},{\bm{v}}_{0}$) is a solution of the Euler and continuity equations for these variables. The dynamics of the background field $\delta\theta_{1}$ is governed by the nonlinear equation (53), which here takes the somewhat simplified but still rather involved form
$\begin{split}\partial_{t}\Big{[}-\frac{1}{\lambda(t)}\big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\big{)}\Big{]}+\nabla\cdot\Big{[}\rho_{0}\nabla\delta\theta_{1}-\frac{\epsilon}{\lambda(t)}(\nabla\delta\theta_{1})\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\Big{]}\\ +\frac{1}{\lambda(t)}\left.\frac{\partial f_{1}}{\partial\rho_{c}}\right|_{0}\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}-\left.\frac{\partial f_{1}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}-\frac{\epsilon}{\lambda(t)}\left.\frac{\partial^{2}f_{1}}{\partial\rho_{c}\partial\theta_{c}}\right|_{0}\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\delta\theta_{1}=0.\end{split}$ (59)
Solving this nonlinear partial differential equation for a given $U(t)$, even for the simplest case of instantaneously switching from zero to a finite $U$, where it simplifies considerably, and proving the convergence of the solution is a nontrivial task. We defer this to a future study. Solving similar classes of partial differential equations is performed for example by the Poincaré-Lighthill-Kuo method [50]. We can however conclude already from the level of complexity of Eq. (59) that the renormalized metric will show features not present in the bare FLRW metric.
At late times, the perturbative expansion breaks down, and only a fully nonlinear treatment similar to the classical analysis of Ref. [41] will thereafter be applicable. Finally, a shock will be encountered, and dispersive effects, coming in a BEC from density gradients (the quantum pressure which we neglected), need in addition to be considered [52].
VI Conclusion
The intrinsic nonlinearity of every interacting quantum field theory (classically and in the hydrodynamical limit a fluid-dynamical system with a given equation of state) requires in general that one goes beyond the standard linearized Unruh paradigm when formulating an analogue gravity scenario. Specifically, to account self-consistently for effects which stem from the nonlinearity, to second (or higher) order in the fluctuation amplitude, one needs to take backreaction of the fluctuations onto the condensate into account. We have demonstrated in our approach to the quantum many-body problem that this engenders, when enforcing Lorentz invariance of the wave equation for the second-order fluctuations, to leading order in the many-body expansion in powers of $N^{1-/2}$, corrections to the Unruh paradigm metric (which is by definition seen by the first-order fluctuations). To higher order in $N^{-1/2}$, quantum nonlinearity, in addition, creates emergent vector fields. We have furthermore shown that number conservation implies the potential presence of a imaginary part of the frequency for excitations above the background-II for second-order perturbations.
While probing these correcting nonlinear effects on the standard analogue gravity Unruh paradigm will be experimentally challenging, they might in future experiments become testable. For example by increasing, in the highly controlled way feasible in the quantum optical context, the contact coupling $U$ via Feshbach resonances to simulate quantum many-body effects on a FLRW background as discussed above. This, then, takes the detail to which quantum field dynamics in curved spacetime, and thus the background field concept itself, can be tested to another level of accuracy. In experimental terms, this is contingent on the ability to measure higher-order correlation functions than those of the (standard) density-density variety, as experimentally demonstrated, e.g., in Ref. [53].
VII ACKNOWLEDGMENTS
We thank Satadal Datta and Caio César Holanda Ribeiro for helpful discussions. KP would like to acknowledge Arun Rana and Kuntal Pal for instructive conversations. This work has been supported by the National Research Foundation of Korea under Grants No. 2017R1A2A2A05001422 and No. 2020R1A2C2008103.
Appendix
We show here how the quantum backreaction term $\phi_{2}$ affects the evolution of a quantum-backreacted condensate. According to the fluid dynamical description of the BEC, which we have employed here, the primary two equations governing the dynamics are (21) and (22). These two equations in terms of the classical fluid variables are different from the continuity and Euler equations obtained from the GPE (5) due to the presence of $f_{1}$ and $f_{3}$, which encapsulate the backreaction contribution from the depletion cloud. Here, our primary goal is an estimation of the size and development of these contributions in a realistic model of a BEC which accounts for backreaction. We will to this end consider the model studied previously in [36], where an initially noninteracting Bose gas of uniform density at rest, trapped in a hard-walled one-dimensional container (i.e., in a one-dimensional box trap with no lateral movement permitted by the trapping geometry) of total length $\ell$ extending from $x=-\ell/2$ to $x=+\ell/2$, is subjected to an interaction quench at $t=0$. For $t<0$, the gas is described as solution of the GPE $\phi_{0}$ such that $\rho_{0}=|\phi_{0}|^{2}=$ constant and is at rest. In the noninteracting regime ($U=0$), the density of the depleted cloud $\rho_{\hat{\chi}}=0$, and consequently backreaction is also not present, i.e., $\phi_{2}=0$. For $t\geq 0$, the leading-order condensate order parameter $\phi_{0}$ remains unchanged, and quantization of the field $\hat{\chi}$, along with the condensate-correction described by $\phi_{2}$, describes the coupled evolution of the condensate and non-condensate parts fully analytically [38, 37].
To compute the function $f_{1}(x,t)$ related to the real part of the anomalous correlator, we resort to the continuity equation satisfied by the background $\phi_{c}$, Eq. (18), to obtain
$f_{1}(x,t)=\partial_{t}\rho_{\phi_{c}}+\partial_{x}J_{\phi_{c}}.$ (60)
We then use that for the solution of Ref. [38] we have $\rho_{\phi_{c}}=\rho_{\phi_{0}}+\rho_{\phi_{2}}+\mathcal{O}(N^{-1/2})$ and $J_{\rm c}=J_{\phi_{0}}+J_{\phi_{2}}+\mathcal{O}(N^{-1/2})$, where $J_{\phi_{0}}=\mbox{Im}[\phi^{*}_{0}\partial_{x}\phi_{0}]$ and
$\rho_{\phi_{2}}$ $=2\mbox{Re}\ [\phi_{0}^{*}\phi_{2}],$ (61)
$J_{\phi_{2}}$ $=\mbox{Im}\ [\phi_{0}^{*}\partial_{x}\phi_{2}+(\partial_{x}\phi_{0})\phi_{2}^{*}].$ (62)
Since the analytical solution for $\phi_{c}$ in terms of $\phi_{0}$ and $\phi_{2}$ is known, it is possible to compute $f_{1}(x,t)$ from (60), setting $\phi_{0}=\exp(-i\mu_{0}t)\sqrt{\rho_{0}}$, where $\mu_{0}=U\rho_{0}$ is the (lowest order) chemical potential, and the lowest-order density $\rho_{0}$ is constant, which defines the bulk coherence length scale via $1/\xi_{0}^{2}=U\rho_{0}=c_{s0}^{2}$, so that all quantities are expressed in powers of $\xi_{0}$.
For being self-contained, we quote here in full the form of the second-order contributions to density $\rho_{\phi_{2}}$ and current $J_{\phi_{2}}$, for $t\geq 0$ (where $t$ is in units of $\xi_{0}^{2}$), from Ref. [36]
$\rho_{\phi_{2}}$ $=$ $-\frac{t^{2}}{\ell}-\frac{1}{4\ell}\sum_{n=1}^{\infty}\frac{(-1)^{n}}{\omega_{n}^{2}}\left\{2(-1)^{n}[1-\cos(2\omega_{n}t)]+\cos(2k_{n}x)\left[\frac{2-k_{n}^{2}}{1+k_{n}^{2}}+2\cos(2\omega_{n}t)-\frac{k_{n}^{2}+4}{k_{n}^{2}+1}\cos(\omega_{2n}t)\right]\right\}\!,$ (63)
$J_{\phi_{2}}$ $=$ $\frac{-2}{\ell}\sum_{n=1}^{\infty}\frac{(-1)^{n}\sin(2k_{n}x)}{k_{n}}\left[\frac{\sin(2\omega_{n}t)}{2\omega_{n}}-\frac{\sin(\omega_{2n}t)}{\omega_{2n}}\right].$ (64)
These expressions can be obtained from the backreaction field $\phi_{2}$, which was solved for exactly in [36]. Here, we have indicated mode frequencies $\omega_{n}=\sqrt{k_{n}^{2}(k_{n}^{2}/4+1)}$, and wavevectors $k_{n}=n\pi/\ell$, which are a result of imposing Neumann boundary conditions at $x=\pm\ell/2$.
We present in Figs. A.1 (a) and (b) selected plots of the variation of $f_{1}(x,t)$ with $x$ at various instants $t\geq 0$. We note that for all profiles the function $f_{1}<0$, which confirms that there is an outgoing particle flux leaving the condensate (particle loss by the condensate). After the system is quenched from a noninteracting to an interacting regime, $f_{1}$ becomes nonzero, following a pattern similar to that of $\rho_{\hat{\chi}}$ [36], and such that the growth in the bulk of the system is not sensitive to the boundary region.
We also plot in Fig. A.2 the evolution of $f_{1}$ with time for different locations inside the trap. We infer that the growth of $f_{1}$ at different $x$, just after the interaction is turned on, follows the same universal pattern. It is dominated by a quadratic growth, which is the contribution of the zero mode contained in (63). This growth is universal irrespective of the location inside the trap. Furthermore, the evolution is drastically different depending on the location in the trap. Near the center of the gas, it is oscillatory. However, close to the hard wall, the contribution to the backreaction quickly saturates with time. These observations are consistent with the observations made in Ref. [37], that as the interaction is suddenly turned on, the energy exchange of the condensate with the quasiparticle excitation cloud initially occurs very fast, and then saturates to an oscillatory behavior, within a time scale depending on the position in the trap.
Figure A.1: (a) Temporal evolution of the spatial dependence of $f_{1}$ for a condensate of total length $\ell=40$, extending from $x=-20$ to $x=20$. Red: $t=0.5$, Black: $t=1$, Blue: $t=2$, Green: $t=3$, Magenta: $t=4$. (b) Evolution of $f_{1}$ for a condensate of larger size $\ell=100$; instants of time indicated with identical colors as in (a). We have verified that the plots do not visibly differ for mode numbers $n\gtrsim 300$; $x$ is in units of $\xi_{0}$ and $f_{1}$ is in units of $1/\xi_{0}^{3}$.
Figure A.2: Temporal evolution of $f_{1}$ for a condensate of total length $\ell=40$ at several locations $x$, parameters as in Fig. A.1 (a). Red: $x=5$, Black: $x=10$, Blue: $x=15$, Green: $x=17$, Magenta: $x=19$. Note that just after the interaction is turned on, there is a universal growth of particle loss from the condensate, irrespective of the position in the trap; $x$ is in units of $\xi_{0}$ and $f_{1}$ is in units of $1/\xi_{0}^{3}$.